Hello,
Please help me with this story problem...I am terrible at them. I got the answer 1.5 ft. but want to check.
A rug is to fit in a room so that a border of even width is left on all four sides. If the room is 12 ft by 15ft and the area of the rug is 108 ft, how wide will the border be? Thanks..>Tammy

Answer by edjones(8007)

room area = 12*15=180 SQ. ft. (don't forget the word "sq.")
rug area = 108 SQ. ft. (don't forget the word "sq.")
Let x=amount of space between wall and rug.
(12-x)(15-x)=108
180-27x+x2=108
rearange and subtract 108 from each side: x^2-27x+72=0
Factor:(x-24)(x-3)=0
x=24 and x=3
x=3 ft. must be the answer because x=24 is WAY too big.
Check:
12-3 * 15-3=9*12=108
Ed